Iron metal is produced in a blast furnace by the reaction of iron (lll) oxide and a coke (pure carbon). If 25.0 moles of pure Fe2O3 is used, how many grams of iron can be produced? The balanced chemical equation for the reaction is: Fe2O3 + 3C --> 2Fe + 3CO
1
Expert's answer
2013-01-21T10:56:04-0500
25 n
Fe2O3 + 3C --> 2Fe + 3CO
& 1 2

Fe2O3 can produce n mol of Fe

n =50 mol

Now you can calculate mass of Fe ( Mw of Fe =55.85)

n=m/Mw

m=n*Mw
m=50*55.85=2794 g